2-(2/3-X)+1/6=-x+5/9 We are supposed to find the solution. Please help me with the steps. There are several more problems like this on my work.
Answer by stanbon(75887)

2-(2/3-X)+1/6=-x+5/9
--------
2-(2/3)+x+(1/6) = -x + (5/9)
(4/3) + x + (1/6) = -x +(5/9)
9/6 +2x = 5/9
2x = (5/9)-(9/6)
2x = (5/9)-(3/2)
2x = -17/18
x = -17/36
--------
Cheers,
Stan H.